25 June 2000  Page 1 of 1  J3/00-241

Subject: Issue 267 From: Van Snyder

1 Introduction

Issue 267 is concerned with the question whether an entire structure is required to be defined if it is an operand.

2 Edits

Edits refer to 00-007r2. Page and line numbers are displayed in the margin. Absent other instructions, a page and line number or line number range implies all of the indicated text is to be replaced by immediately following text, while a page and line number followed by + indicates that immediately following text is to be inserted after the indicated line. Remarks for the editor are noted in the margin, or appear between [ and ] in the text.

[Editor: After “When” insert “an ultimate component of”. Replace “all of the components” 120:26 by “either directly or as a consequence of reference to the structure, it”.]

[Editor: Delete “The 3rd para ... writing this up” – the first three paragraphs and the first 121:3-17 sentence of the fourth paragraph – of issue 267.]